OpenStudy (anonymous):

Find the solution of the higher order differential y'''' - y'' + y' - y = 0

OpenStudy (lgbasallote):

are you familiar with something called "auxilliary solution" or something like that?

OpenStudy (anonymous):

Not at the moment.

OpenStudy (lgbasallote):

auxilliary solution is something like replacing the differential operator with a variable for example: y''' - y'' + y' - y = 0 becomes m^3 - m^2 + m - 1 = 0 does that look familiar?

OpenStudy (anonymous):

Oh I made a typo, I mean y''' - y '' + y' - y, the highest power was supposed to be 3. I know the the characteristic equation is r^3-r^2+r -1

OpenStudy (lgbasallote):

yes that's the auxiliary solution i was referring to

OpenStudy (lgbasallote):

r^3 - r^2 + r - 1 = 0 do you know how to factor it out?

OpenStudy (anonymous):

I'm kind of brain dead at the moment, kind of forgot how to factor out higher order polynomials :( It's r^2(r-1)+(r-1)=0 right? So we have a double root at 1, and a real root at 0?

OpenStudy (lgbasallote):

no...you still have to factor r^2(r-1) + (r-1) further

OpenStudy (lgbasallote):

here's a hint: factor out (r-1) because it's a common factor to both terms...what do you get?

OpenStudy (anonymous):

Blaaaahh, can't believe I grossly overlooked that. (r^2+1), r-1

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so your roots are \[r = 1\] and \[r = \pm i\]

OpenStudy (anonymous):

right so the relating equation would be C1E^t + C2 cos t + C3 sin t ?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

i just don't know why you used t

OpenStudy (anonymous):

Oh I got used to my professor using t... haha

OpenStudy (lgbasallote):

anyway..that would be the right answer

OpenStudy (anonymous):

Thank you.